# Optimal CO2 storage management considering safety constraints in multi-stakeholder multi-site CCS projects: a game theoretic perspective

### v1

@@ 2.3.2 Multi-agent reinforcement learning formulation
To address the inherent interdependencies and competing objectives in multi-stakeholder geological carbon storage (GCS), we reformulate the optimization problem as a constrained Markov game as illustrated in Figure 4. In this framework, each operator is represented as an autonomous agent that interacts with the shared reservoir environment while pursuing its individual economic goals subject to safety constraints. The goal of each agent is to maximize its expected discounted cumulative reward:
$J_{i}(\pi_{1},\dots,\pi_{n})=\mathbb{E}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}r^{i}_{t}\Bigg{]},$
where $\pi_{i}$ denotes the policy of agent $i$.
@@
Each agent $i$ is associated with a cost/penalty function $c^{i}_{t}$, representing violations of operational safety thresholds (e.g., pore pressure exceeding allowable threshold pressure). The optimization objective thus becomes:
$\max_{\pi_{i}}J_{i}(\pi)=\mathbb{E}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}r^{i}_{t}\Bigg{]},\quad\text{s.t.}\quad\mathbb{E}\Bigg{[}\sum_{t=0}^{\infty}\gamma^{t}c^{i}_{t}\Bigg{]}\leq d_{i},\;\;\forall i,$
where $d_{i}$ denotes the maximum allowable cumulative violation for agent $i$. More details of this formulation has been presented in section 3.3.
3 Preliminaries
@@ 3.1 Deep Reinforcement Learning (DRL)
Deep Reinforcement Learning (DRL) is a branch of machine learning designed for sequential decision-making problems where labeled training data are unavailable and agents must learn optimal behaviors through repeated interactions with the environment. DRL has registered tremendous success in solving various sequential decision-making problems, including but not limited to: robotics control and maneuver (Kober et al., 2013; Li et al., 2024; Yu et al., 2025a), mechanical system design (Chen et al., 2025a), large language models (Havrilla et al., ) and so on. A DRL framework is typically formalized as a Markov Decision Process (MDP) (Figure 3 (a)), in which an agent observes the current state $s$, selects an action $a$, and receives a corresponding reward $r$. The objective is to learn a policy that maximizes the cumulative expected reward over a defined time horizon.
@@
$J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right]$ (9)
@@ 3.2 Multi-Agent Reinforcement Learning (MARL)
When multiple agents interact within the same environment, it becomes a Multi-Agent Reinforcement Learning (MARL) problem (Tampuu et al., 2017; Canese et al., 2021), which is often formulated as a Markov game as in Figure 3 (b) (Wang et al., 2022). In MARL, each agent receives observations $o_{i}$, takes actions $a_{i}$, and obtains rewards $r_{i}$, while the environment evolves according to the joint actions $a_{1},...,a_{i},...a_{n}]$ of all agents. Agents may cooperate, compete, or operate under mixed motives. Cooperation aims to maximize a shared reward, while competitive agents pursue individual objectives, particularly zero-sum objectives. This structure makes MARL well-suited for modeling multi-stakeholder carbon storage projects, where operators, landowners, and regulators pursue distinct but interconnected goals.
@@
Figure 3: Schematic diagrams of (a) Markov decision process and (b) Markov game, which correspond to the frameworks for (a) single- and (b) multi-agent RL, respectively. Adapted from (Zhang et al., 2021)
@@
$\langle n,\mathcal{S},\{\mathcal{A}_{i}\}_{i=1}^{n},\mathbb{T},\{r_{i}\}_{i=1}^{n},\gamma\rangle,$
@@
    an action space $\mathcal{A}_{i}$,
@@
    a reward function $r_{i}(s,a_{1},\ldots,a_{n})$,
@@
    a policy $\pi_{i}(a_{i}\mid s)$.
The environment transitions according to the joint action $a=(a_{1},\ldots,a_{n})$:
$s_{t+1}\sim\mathbb{T}(\cdot\mid s_{t},a_{1}^{t},\ldots,a_{n}^{t}).$ (10)
In a multi-stakeholder CCS setting, the transition function should comply with subsurface dynamics and can be approximated by a surrogate model as in equation 4. The objective of each agent in MARL framework is to maximize its own expected return:
$J_{i}(\pi_{1},\ldots,\pi_{n})=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}\right]$ (11)
@@ 3.3 Multi-Agent Reinforcement Learning (MARL) with Safety Constraints
Conventional MARL does not explicitly account for safety, which is critical in domains where safety is of paramount importance. To address this, safe reinforcement learning (Achiam et al., 2017; Gu et al., 2024; Zheng and Gu, 2024) extends MARL by integrating safety constraints into the decision-making process (Figure 4). Each agent receives not only rewards $r_{i}$ but also penalties or costs $c_{i}$ when actions or resulting system states violate predefined safety thresholds. This constrained Markov game (Altman, 2021) formulation restricts agents to a feasible action space that ensures compliance with operational and regulatory safety requirements – in the case of CO2 storage, limiting reservoir pressure to prevent caprock fracturing. By embedding safety directly into the learning process, this framework ensures that optimal strategies are not only economically beneficial but also operationally viable and environmentally secure.
@@
A safe multi-agent reinforcement learning framework tailored for GCS is displayed in 5. Each operator/agent is interacting with the subsurface environment and receiving reward $r_{i}$ and cost $c_{i}$ at each timestep. Based on the observation $o_{i}$, each operator chooses their own well control strategy $a_{i}$ and the joint actions of all operators $a_{1}$, $a_{i}$, …, $a_{n}$ drive the current environment state $s^{t}$ to next environment state $s^{t+1}$.
@@
$\langle n,\mathcal{S},\{\mathcal{A}_{i}\}_{i=1}^{n},\mathbb{T},\{r_{i}\}_{i=1}^{n},\{c_{i}\}_{i=1}^{n},\gamma\rangle,$
where each agent additionally has a cost function $c_{i}(s,a)$ representing safety violations (e.g., exceeding reservoir pressure limits). Each agent now optimizes:
$\max_{\pi_{i}}\;J_{i}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{i}^{t}\right]$ (12)
subject to the safety constraint:
$\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{i}^{t}\right]\leq d_{i},\quad\forall i$ (13)
where $d_{i}$ is the allowable safety threshold for agent $i$.
@@
$\mathcal{L}_{i}(\pi,\lambda_{i})=J_{i}(\pi)+\lambda_{i}\left(\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{i}^{t}\right]-d_{i}\right),$ (14)
with $\lambda_{i}\geq 0$ being the Lagrange multiplier balancing reward maximization and constraint enforcement. This coefficient is dynamically updated to penalize constraint violations.
4 Methodologies
@@ 4.1 Safe Multi-Agent DDPG (MADDPG) for GCS
To ensure each operator injects safely while maximizing potentially competing objectives in basin-scale CCS, we adopt the Multi-Agent Deep Deterministic Policy Gradient (MADDPG) framework (Lowe et al., 2017). This framework employs centralized training with decentralized execution to enable multiple operators—modeled as agents—to learn coordinated, well-control policies.
In this setup, each agent $i$ has its own actor–critic network. The deterministic actor $\mu_{\theta_{i}}(o_{i})$ maps the agent’s local observation (e.g., pressures, plume size) to continuous control actions (e.g., injection/extraction rates, bottom-hole pressures). During training, a centralized critic $Q^{r}_{i}(s,a_{1},\dots,a_{N})$ evaluates the joint state–action space, leveraging observations $o_{i}$ and actions $a_{i}$ from all agents to mitigate non-stationarity and provide informative gradients. In execution, however, each agent relies solely on its actor network and local observations, maintaining a fully decentralized decision-making process.
To enforce safety constraints, we extend MADDPG with Lagrangian relaxation. Alongside the reward critic $Q_{i}^{r}$, a cost critic $Q_{i}^{c}$ estimates expected constraint violations (e.g., grid-cell pressures exceeding threshold pressure). The actor for agent $i$ is trained using the Lagrangian objective:
$\mathcal{L}(\pi_{i})\;=\;\,\mathbb{E}\Big{[}-\,Q_{i}^{r}(s,a)\;+\;\lambda_{i}\,Q_{i}^{c}(s,a)\,\Big{]},$ (15)
where $\lambda_{i}\geq 0$ is a non-negative weight that controls how strongly safety violations are penalized. It is automatically adjusted during training—if an agent is breaking the safety limit too often, $\lambda_{i}$ increases to apply more penalty. The critics are learned through temporal-difference (TD) regression with experience replay and target networks; the actors are updated via deterministic policy gradients; and target networks are soft-updated to stabilize learning. The detailed training and execution scheme was visualized in 15 and the algorithm is shown in 1. This approach yields policies that maximize economic return while ensuring pressure safety, aligning directly with our constrained Markov game formulation for multi-stakeholder CCS.
Figure 5: A safe multi-agent reinforcement learning framework tailored for multi-stakeholder CCS projects. Each agent/operator is interacting with subsurface dynamical environment, receiving immediate reward (defined by present values) and penalty (proportional to number of gridblocks that exceed the threshold pressure) and driving the environment to next state.
@@ 4.2 Adaptations for GCS Optimization
    Constraint-aware learning: Safety constraints are integrated via cost critics and Lagrangian multipliers, ensuring that pore pressure at well locations or each grid cell remains below the allowable threshold pressure ($p_{\text{grid, i}}\leq\alpha p_{\text{frac, i}}$).
4.3 Reward and Penalty Modeling for Geological CO2 storage
In reinforcement learning, a reward model defines the feedback an agent receives for its actions, signaling which behaviors are beneficial and guiding the agent toward desired outcomes (Yu et al., 2025b; Arora and Doshi, 2021). In safe reinforcement learning, this is complemented by a penalty model, which assigns negative feedback when actions violate safety constraints. Together, reward and penalty models ensure that agents not only maximize designated performance but also operate within required safety and risk boundaries. In multi-agent reinforcement learning (MARL) for geological CO2 storage (GCS), rewards and penalties must capture the trade-offs between economic returns and regulatory compliance, while accounting for the dynamic, time-dependent nature of subsurface processes.

### v2

@@ 2.3.2 Constrained Markov game formulation
To address the inherent interdependencies and competing objectives in multi-stakeholder geological carbon storage (GCS), we reformulate the optimization problem as a constrained Markov game as illustrated in Figure 4. Specifically, we apply multi-agent reinforcement learning (MARL) to learn an optimal policy for each stakeholder operating in the shared reservoir environment. In this framework, each operator is represented as an autonomous agent that interacts with the shared reservoir environment while pursuing its individual economic goals subject to safety constraints. The goal of each agent is to maximize its long-term return, defined as:
$\max_{\pi^{i}}\;J_{i}\!\left(\pi^{i},\pi^{-i}\right)=\mathbb{E}\!\left[\sum_{t=0}^{T}\gamma_{t}r_{t}^{i}\;\middle|\;s_{0},\;a_{t}^{i}\sim\pi^{i}(\cdot\mid s_{t}),\;a_{t}^{-i}\sim\pi^{-i}(\cdot\mid s_{t})\right]$ (10)
where $\pi^{i}$ denotes the policy of agent $i$, $\pi^{-i}$ represents other agents’ policies except agent $i$, $\gamma_{t}$ is the discount factor, $r_{t}^{i}$ is the immediate reward (PV) received by agent $i$ through interacting with the reservoir environment, and $a_{t}^{i}$ represents the actions from operator/agent i, mapped from the respective policies conditioned on the current state $s_{t}$. More details of this formulation have been presented in section 3.2.
@@
In constrained Markov game, each agent $i$ is associated with a cost/penalty function $c^{i}_{t}$, representing violations of operational safety thresholds (e.g., pore pressure exceeding allowable threshold pressure). Analogous to the reward $r_{t}^{i}$, this penalty $c^{i}_{t}$ depends on current state $s_{t}$, action $a_{t}$ and subsequent state $s_{t+1}$. Collectively, the objective of this constrained Markov game formulation thus becomes:
$\max_{\pi^{i}}$ $J_{i}\!\left(\pi^{i},\pi^{-i}\right)$ (11)
s.t. $\mathbb{E}\!\left[\sum_{t=0}^{T}\gamma^{t}c_{t}^{i}\;\middle|\;s_{0},\;a_{t}^{i}\sim\pi^{i}(\cdot\mid s_{t}),\;a_{t}^{-i}\sim\pi^{-i}(\cdot\mid s_{t})\right]\leq d^{i},$
where $d^{i}$ denotes the maximum allowable cumulative violation for agent $i$.
The constrained Markov game paradigm enables real-time, closed-loop decision making in basin-scale carbon storage, where storage operations are sequential, state-dependent. Operators’ actions must adapt to evolving subsurface conditions (e.g., pressure and plume migration) and remain feasible under safety limits. The constrained Markov game directly models inter-operator interactions and learns decentralized policies that map observations/states to actions while enforcing regulatory or safety constraints.
3 Background on Markov Games and Multi-Agent Reinforcement Learning
@@ 3.1 Deep Reinforcement Learning (DRL)
Deep Reinforcement Learning (DRL) is a branch of machine learning designed for sequential decision-making problems where labeled training data are unavailable and agents must learn optimal behaviors through repeated interactions with the environment. DRL has registered tremendous success in solving various sequential decision-making problems, including but not limited to: robotics control and maneuver [31, 33, 56], mechanical system design [14], large language models [24] and so on. A DRL framework is typically formalized as a Markov Decision Process (MDP) (Figure 3 (a)), in which an agent observes the current state $s$, selects an action $a$, and receives a corresponding reward $r$. The objective is to learn a policy that maximizes the cumulative expected reward over a defined time horizon.
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\;\Big|\;a_{t}\sim\pi(\cdot\mid s_{t}),\,s_{0}\right]$ (12)
Accordingly, it is convenient to summarize the long-term consequence of decisions through value functions. The state-value function under policy $\pi$ is defined as the expected discounted return starting from state $s$,
$V_{\pi}(s):=\mathbb{E}\!\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\,\middle|\,s_{0}=s,\;a_{t}\sim\pi(\cdot\mid s_{t})\right],$ (13)
and the action-value function is the expected discounted return starting from $(s,a)$,
$Q_{\pi}(s,a):=\mathbb{E}\!\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\,\middle|\,s_{0}=s,\;a_{0}=a,\;a_{t}\sim\pi(\cdot\mid s_{t})\right].$ (14)
Intuitively, $V_{\pi}(s)$ evaluates how good it is to be in state $s$ under $\pi$, while $Q_{\pi}(s,a)$ evaluates how good it is to take action $a$ at state $s$ and then follow $\pi$, which directly supports action selection and policy improvement. In this context, value-based methods learn $V_{\pi}$ or $Q_{\pi}$ from data and derive a policy by selecting actions that maximize the estimated value (e.g., $\arg\max_{a}Q(s,a)$), whereas policy-based methods parameterize the policy through $\pi_{\theta}$ and optimize the policy parameters $\theta$ directly (typically via policy gradients) without requiring an explicit maximization over actions; many modern actor–critic algorithms combine both by using a learned value function to guide and stabilize policy updates.
@@ 3.2 Multi-Agent Reinforcement Learning (MARL)
Traditional DRL algorithms cannot cope with multiple agents as the environment becomes non-stationary. When multiple agents interact within the same environment, it becomes a Multi-Agent Reinforcement Learning (MARL) problem [46, 10], which is often formulated as a Markov game as in Figure 3 (b) [50]. In MARL, each agent receives observations $o_{i}$, takes actions $a_{i}$, and obtains rewards $r_{i}$, while the environment evolves according to the joint actions $a_{1},...,a_{i},...a_{n}]$ of all agents.
@@
Figure 3: Schematic diagrams of (a) Markov decision process and (b) Markov game, which correspond to the frameworks for (a) single- and (b) multi-agent RL, respectively. Adapted from [58]
@@
$\langle n,\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{n},\mathbb{T},\{r^{i}\}_{i=1}^{n},\gamma\rangle,$
@@
    an action $a^{i}$ from action space $\mathcal{A}^{i}$,
@@
    a reward function $r^{i}(s,a^{1},\ldots,a^{n})$,
@@
    a policy $\pi^{i}(a^{i}\mid s)$.
The environment transitions according to the joint action $a=(a^{1},\ldots,a^{n})$:
$s_{t+1}\sim\mathbb{T}(\cdot\mid s_{t},a_{t}^{1},\ldots,a_{t}^{n}).$ (15)
In a multi-stakeholder CCS setting, the transition function should comply with subsurface dynamics and can be approximated by a surrogate model as in equation 4. The objective of each agent $i$ in MARL framework is to maximize its own long-term return by finding a state-conditioned policy $\pi^{i}(\cdot\mid s)$ while considering joint policy $\pi^{-i}(\cdot\mid s)$ of all other agents except $i$, as presented in equation 10.
Similarly, the agent-specific state-value functions and action-value functions can be defined as follows:
$V^{i}(s):=\mathbb{E}\!\left[\sum_{t=0}^{T}\gamma^{t}r_{t}^{i}\;\middle|\;s_{0}=s,\;a_{t}^{i}\sim\pi^{i}(\cdot\mid s_{t}),\;a_{t}^{-i}\sim\pi^{-i}(\cdot\mid s_{t})\right],$ (16)
$Q_{r}^{i}(s,a^{i},a^{-i}):=\mathbb{E}\!\left[\sum_{t=0}^{T}\gamma^{t}r_{t}^{i}\;\middle|\;s_{0}=s,\;a_{0}^{i}=a^{i},\;a_{0}^{-i}=a^{-i},\;a_{t}^{i}\sim\pi^{i}(\cdot\mid s_{t}),\;a_{t}^{-i}\sim\pi^{-i}(\cdot\mid s_{t})\right].$ (17)
Agents may cooperate, compete, or operate under mixed motives. Cooperation aims to maximize a shared reward, while competitive agents maximize their own objectives. In general, such interactions are general-sum, where there is no restrictions imposed on the goals of agents and their rewards can be arbitrarily related [26, 45]. The competitive CCS scenario considered in this work is in the general-sum sense, since each operator maximizes its own NPV and the sum of total NPV is not fixed. This structure makes MARL well-suited for modeling multi-stakeholder carbon storage projects, where operators, landowners, and regulators pursue distinct but interconnected goals.
@@ 3.3 Multi-Agent Reinforcement Learning (MARL) with Safety Constraints
Conventional MARL does not explicitly account for safety, which is critical in domains where safety is of paramount importance. To address this, safe reinforcement learning [1, 21, 59] extends MARL by integrating safety constraints into the decision-making process (Figure 4). Each agent receives not only rewards $r^{i}$ but also penalties or costs $c^{i}$ when actions or resulting system states violate predefined safety thresholds. This constrained Markov game [2] formulation restricts agents to a feasible action space that ensures compliance with operational and regulatory safety requirements – in the case of CO2 storage, limiting reservoir pressure to prevent caprock fracturing. By embedding safety directly into the learning process, this framework ensures that optimal strategies are not only economically beneficial but also operationally viable and environmentally secure.
@@
A safe multi-agent reinforcement learning framework tailored for GCS is displayed in figure 5. Each operator/agent is interacting with the subsurface environment and receiving reward $r^{i}$ and cost $c^{i}$ at each timestep. Based on the observation $o^{i}$, each operator chooses their own well control strategy $a^{i}$ and the joint actions of all operators $a^{1}$, $a^{i}$, …, $a_{n}$ drive the current environment state $s_{t}$ to next environment state $s_{t+1}$. In field operations, injection controls are typically updated at discrete management intervals rather than continuously. Accordingly, in our formulation each action $a_{t}$ represents a piecewise-constant well-control decision applied over a user-defined control period $\Delta t$. This abstraction is intended to reflect practical operational scheduling while retaining a sequential decision-making structure suitable for MDP/Markov-game formulations.
@@
$\langle n,\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{n},\mathbb{T},\{r^{i}\}_{i=1}^{n},\{c^{i}\}_{i=1}^{n},\gamma\rangle,$
where each agent additionally has a cost function $c^{i}(s,a)$ representing safety violations (e.g., exceeding reservoir pressure limits). Building upon this framework, the problem can be formulated as in equation 11. In similar vein,
$Q_{c}^{i}(s,a^{i},a^{-i}):=\mathbb{E}\!\left[\sum_{t=0}^{T}\gamma^{t}c_{t}^{i}\;\middle|\;s_{0}=s,\;a_{0}^{i}=a^{i},\;a_{0}^{-i}=a^{-i},\;a_{t}^{i}\sim\pi^{i}(\cdot\mid s_{t}),\;a_{t}^{-i}\sim\pi^{-i}(\cdot\mid s_{t})\right].$ (18)
estimates the expected constraints violations.
@@
$\mathcal{L}^{i}(\pi,\lambda^{i})=J^{i}(\pi)+\lambda^{i}\left(\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{t}c^{i}_{t}\right]-d^{i}\right),$ (19)
with $\lambda^{i}\geq 0$ being the Lagrange multiplier balancing reward maximization and constraint enforcement. This coefficient is dynamically updated to penalize constraint violations.
4 Proposed Framework - Constrained Markov Game for GCS
@@ 4.1 Safe Multi-Agent DDPG (MADDPG) for GCS
To ensure each operator injects safely while maximizing potentially competing objectives in basin-scale CCS, we adopt the Multi-Agent Deep Deterministic Policy Gradient (MADDPG) framework [35]. This framework employs centralized training with decentralized execution (CTDE) to enable multiple operators—modeled as agents—to learn coordinated, well-control policies.
In this setup, the centralized critic network $Q_{r}^{i}(s,a^{i},a^{-i})$ is used only during training to evaluate action values under the joint state–action information and to provide a stable learning signal; it does not determine or execute actions. Each agent’s actor network $\mu_{\theta^{i}}(o^{i})$ outputs its own action based on local observations (e.g., pressures, plume saturation) . The economic objective optimized by each agent depends on the coalition structure through the reward formulation: in the fully cooperative case, agents share a common team reward (total NPV), whereas in the competitive (or mixed) cases each agent maximizes its own expected discounted economic return. The complete set of coalition structures is summarized in Table 1.
In this study, each operator’s observation is restricted to an operator-centric set (e.g., bottom-hole pressure and locally available variables) to reflect limited data availability in GCS settings. Pressure interference from neighboring projects is still captured through the shared reservoir dynamics and is reflected in each operator’s measurements and constraint costs. If boundary or regional pressure monitoring is available or mandated (e.g., observation wells near permit boundaries or regulator-provided measurements), these data can be incorporated directly into the observation vector $o_{i}$ as additional channels without changing the proposed formulation.
Figure 5: A safe multi-agent reinforcement learning framework tailored for multi-stakeholder CCS projects. Each agent/operator is interacting with subsurface dynamical environment, receiving immediate reward (defined by present values) and penalty (e.g. injection pressure exceeding the fracturing pressure) and driving the environment to next state.
@@ 4.2 Adaptations for GCS Optimization
    Compliance-aware learning: Safety is enforced by incorporating a cost critic network $Q^{i}_{c}$ and a Lagrangian multiplier $\lambda^{i}$, ensuring that pore pressure at well locations or each grid cell remains below the allowable threshold pressure ($p_{\text{grid}}\leq\alpha p_{\text{frac}}$). Specifically, the actor network for agent $i$ is trained using the Lagrangian objective:
    $\mathbb{E}\Big[-\,Q_{r}^{i}(s,a)\;+\;\lambda^{i}\,Q_{c}^{i}(s,a)\,\Big],$ (20)
    where $\lambda^{i}\geq 0$ is a non-negative weight that controls how strongly safety violations are penalized. It is automatically adjusted during training—if an agent is breaking the safety limit too often, $\lambda^{i}$ increases to apply more penalty. The critics are learned through temporal-difference (TD) regression with experience replay and target networks; the actors are updated via deterministic policy gradients; and target networks are soft-updated to stabilize learning. The detailed training and execution scheme was displayed in Figure 13 and the algorithm is shown in Algorithm 1.
It should be noted that the pressure limit used in this work is an illustrative and widely adopted safety constraint, but it may not be the binding constraint in every basin-scale setting (e.g., highly open boundaries or low injection intensity). The proposed constrained Markov game/safe-MARL framework is not restricted to pressure: other operational or regulatory constraints (e.g., plume extent/area of review (AoR), leakage-risk indicators) can be incorporated by defining the agent-specific cost function $c_{i}(s,a)$ accordingly. Nevertheless, even at basin scale, pressure can become limiting when multiple operators inject concurrently and interference accumulates over time.
4.3 Reward and Penalty Formulation for Geological CO2 Storage
In reinforcement learning, a reward model defines the feedback an agent receives for its actions, signaling which behaviors are beneficial and guiding the agent toward desired outcomes [57, 4]. In safe reinforcement learning, this is complemented by a penalty model, which assigns negative feedback when actions violate safety constraints. Together, reward and penalty models ensure that agents not only maximize designated performance but also operate within required safety and risk boundaries. In multi-agent reinforcement learning (MARL) for geological CO2 storage (GCS), rewards and penalties must capture the trade-offs between economic returns and regulatory compliance, while accounting for the dynamic, time-dependent nature of subsurface processes.